Solve this differential equations by finding the coefficientes A0, An and bn for the fourier series

Encuentre los Coes o,n y bn Para ka seriede Fourier S Fte 10 )0XS10
Answer #1

S)=10(-1) de hae, Trean The Fouhior seriv is Giten +mings Sin3nTk a Cas nTY where 10 o- + Sin nT an Cas nT LO uhere t n = -onT d -1) o Sin nT (n-1) an An=2 h T lo Sin2n-o +2 an= CanITY n IT -H hT dn 2xio CrnT -D 4o n T daly Suin nr dw n lo C nn 11/5l0 +0 + 5 n InT hIT n = 3 3 n Hence Se OD uantaas